solve by substitution
x+y=-9
x-y=14

Answer by checkley71(8403)

X+Y=-9 OR X=-9-Y THEN -9-Y-Y=14 OR -9-2Y=14 OR -2Y=14+9 OR -2Y=23 Y=23/-2 OR Y=-11.5 THEN
X-11.5=-9 OR X=-9+11.5 OR X=2.5